Mercury Inc. purchased equipment in 2019 at a cost of $400,000. The equipment was expected to produce 700,000 units over the next five years and have a residual value of $50,000. The equipment was sold for $210,000 part way through 2021. Actual production in each year was: 2019 = 100,000 units; 2020 = 160,000 units; 2021 = 80,000 units. Mercury uses units-of-production depreciation, and all depreciation has been recorded through the disposal date.
Required:
1. Calculate the gain or loss on the sale.
2. Prepare the journal entry to record the sale.
3. Assuming that the equipment was instead sold for $245,000, calculate the gain or loss on the sale.
4. Prepare the journal entry to record the sale in requirement 3.
Answer #1

Answer 1:

Sale value 210000
Less: Book value:
Cost of equipment 400000
Accumulated depreciation 170000 230000
Gain (Loss) on sale of equipment -20000

Working note:

Calculation of depreciation expense per unit:

Cost = 400000

Residual value = 50000

Depreciable value (400000-50000) = 350000

Expected production = 700000

Depreciation expense per unit = Depreciable value / Expected production

= 350000 / 700000 = 0.50 per unit

Calculation of Accumulated depreciation:

Depreciation expense for 2019 (100000 * 0.50) = 50000

Depreciation expense for 2020 (160000 * 0.50) = 80000

Depreciation expense for 2021 (80000 * 0.50) = 40000

Accumulated depreciation = 170000

Answer 2:

Date Account title and explanation Debit Credit
2021 Cash 210000
Accumulated depreciation 170000
Loss on sale of equipment 20000
Equipment 400000
(to record sale of equipment)

Answer 3:

Sale value 245000
Less: Book value:
Cost of equipment 400000
Accumulated depreciation 170000 230000
Gain (Loss) on sale of equipment 15000

Answer 4:

Date Account title and explanation Debit Credit
2021 Cash 245000
Accumulated depreciation 170000
Equipment 15000
Gain on sale of equipment 400000
(to record sale of equipment)

Dear Student please upvote my answer, Thankyou.